### v1

# From Social Division to Cohesion with AI Message Suggestions in Online Chat Groups

Abstract
Social cohesion is difficult to sustain in societies marked by opinion diversity, particularly in online communication. As large language model (LLM)-driven messaging assistance becomes increasingly embedded in these contexts, it raises critical questions about its societal impact. We present an online experiment with 557 participants who engaged in multi-round discussions on politically controversial topics while freely reconfiguring their discussion groups. In some conditions, participants received real-time message suggestions generated by an LLM, either personalized to the individual or adapted to their group context. We find that subtle shifts in linguistic style during communication, mediated by AI assistance, can scale up to reshape collective structures. While individual-focused assistance leads users to segregate into like-minded groups, relational assistance that incorporates group members’ stances enhances cohesion through more receptive exchanges. These findings demonstrate that AI-mediated communication can support social cohesion in diverse groups, but outcomes critically depend on how personalization is designed.
keywords:
social cohesion, opinion diversity, online communication, large language models, human–AI interaction
1 Introduction
Social cohesion remains a persistent challenge in societies characterized by opinion diversity [1, 2, 3]. Conversations across differences are fraught: people struggle to engage constructively with those who hold opposing views, and stereotypes bias perceptions of outgroup members [4, 5]. The challenges individuals face in conversation are reinforced by structural dynamics, as homophily drives clustering with similar others and limits cross-cutting ties at scale [6, 7, 8]. Online platforms intensify this micro–macro loop by lowering inhibitions in communication, accelerating interpersonal clustering, and amplifying these effects at scale [9, 10, 11, 12, 13]. The consequences—from political polarization to breakdowns in global coordination—underscore the urgency of understanding how opinion diversity is managed within online environments [14, 15, 16, 17, 18].
Recent advances in artificial intelligence (AI) introduce a new layer of influence to this challenge [19, 20]. Large language models (LLMs) are now widely integrated into everyday online communication through messaging assistance tools [21, 22, 23]. While such assistance can reduce the burden of writing [24], conventional personalization features risk amplifying stance differences among people with diverse views [25, 26]. Conversely, some designs for AI-mediated communication have been proposed to foster interpersonal trust and build consensus across lines of disagreement [27, 1, 22, 28, 29]. Yet, it remains unclear how these micro-level interventions can cascade upward to shape macro-level changes in human group composition. How can AI message suggestions foster social cohesion rather than exacerbate division in diverse populations?
Here, we address this question through a preregistered experiment conducted on a customized online chat platform (see Methods for details) [30]. Participants ($N=557$), recruited from Prolific [31], were randomly assigned to sessions of 6 to 15 people ($M=9.42$, $s.d.=2.57$), subdivided into groups of 2 to 4 (Fig. 1). Each group discussed a politically controversial topic selected to maximize opinion diversity for that session, based on participants’ pre-task survey responses (see Extended Data Table 1 for survey items) [27]. Participants were able to update their survey stance during the session, though this occurred rarely ($6.40\%$ of cases).
Participants were instructed to engage in group conversations, but were not asked to reach consensus, reflecting the open-ended and non-teleological nature of informal online discussions [32, 14]. After each three-minute discussion, participants decided whether to remain in their current group, join another, or create a new one. If only one participant remained in a group, they continued the discussion with a chatbot to ensure continuous engagement. This procedure was repeated for four rounds, followed by a final switching decision that determined the Round 5 group composition used for structural analysis.
Within this setup, we manipulated the presence and type of AI assistance in message writing through three experimental conditions [27]. In the no-assistance condition, participants composed all messages themselves and did not receive any AI-generated suggestions. In the other two conditions, participants received real-time message suggestions generated by OpenAI’s GPT-4o model as the group chat unfolded. Although suggestions were tailored to each user based on their survey responses and subsequent updates in both conditions, the personalization context differed between them (see Methods; full prompts appear in SI.).
In the individual assistance condition, the model was prompted to help participants articulate their stance on the conversation topic, aligning suggestions with each individual’s viewpoint. In contrast, in the relational assistance condition, the modelincorporated not only the participant’s stance but also those of their group members, with the goal of fostering mutual understanding. As a result, individual-focused suggestions primarily supported self-expression, whereas relationally aware suggestions more often conveyed receptive language and positive sentiment, even in the presence of disagreement (Extended Data Table 2), as confirmed by quantitative analyses (Extended Data Fig. 1). These suggestions could be sent with a single click, edited, or ignored in favor of self-written messages.
Following the preregistered protocol [30], we excluded sessions with fewer than six participants or with more than 25% dropout. The final dataset includes 20 sessions per condition (60 sessions total), comprising 557 participants overall (187 in the no assistance condition, 182 in the individual assistance condition, and 188 in the relational assistance condition). Each participant completed only one session. Per-session participant counts, dropout rates, and topic assignments did not differ significantly across conditions (Chi-square tests for participant counts: $p=0.946$; dropout rates: $p=0.538$; Fisher’s exact test for topic distributions: $p=0.533$).
We evaluate the effects of AI assistance on group composition and social cohesion, measured through stance assortativity—a session-level indicator of how strongly group composition reflects similarity in participants’ opinions. Individuals are treated as connected if they belong to the same group [33], and assortativity is computed from their stances on the discussion topic (3-3 = totally disagree to +3+3 = totally agree) [34]. Higher assortativity values indicate stronger clustering of like-minded individuals (i.e., greater social division and lower cohesion), whereas lower values indicate more cross-cutting ties (i.e., lower division and greater cohesion) (Fig. 1). Because random assignment produced variation in initial group composition, we analyze changes relative to the first-round baseline. To examine the mechanisms underlying these structural shifts, we also analyze conversational dynamics, including message volume and sentiment, with the latter measured using VADER sentiment analysis [35].
Figure 1: Illustration of group-composition dynamics. Each session begins with 2 to 4 groups, each consisting of 2 to 4 members, shown as gray clusters. Stance assortativity is calculated at the session level from participants’ stances on the discussion topic, represented by node colors. As participants switch groups across rounds, group composition evolves, either increasing social division (positive assortativity) or fostering cohesion (negative assortativity).
2 Results
2.1 Effects of AI Assistance on Social Cohesion
Figure 2: Relational AI assistance reduces stance assortativity. The left panel shows deviations from baseline assortativity (Round 1), and the right panel shows estimated slopes of change with statistical comparisons against the no-assistance condition (Extended Data Table 3). Error bars represent mean $\pm$ s.e.m.
We found that stance assortativity remained stable or increased slightly across rounds in the no-assistance condition, but decreased substantially in the relational assistance condition (Fig. 2). After four rounds, assortativity decreased by an average of 0.135 points on the –1 to 1 scale under relational assistance. This reduction was statistically significant relative to the no-assistance condition ($p=0.004$; regression analysis; Extended Data Table 3). In contrast, changes in assortativity under individual assistance did not differ significantly from the control ($p=0.608$). Relational assistance promoted more cross-cutting group composition, while individual assistance did not alter baseline patterns of clustering.
To further understand the effects of AI assistance on structural dynamics, we examined the number of conversation partners, the number of conversation groups, within-group stance distance, and between-group stance distance across interaction rounds. Participants, on average, increased their conversation partners without AI assistance ($p<0.001$ in the no-assistance condition; Fig. 3a). This trend was observed across all conditions but it was amplified under individual assistance: the increase in conversation partners was significantly larger than in the control ($p=0.028$; Extended Data Table 3). Relational assistance did not differ significantly from the control ($p=0.712$).
The number of groups remained stable across rounds in the control condition ($p=0.280$), and neither AI condition altered this pattern (individual: $p=0.317$; relational: $p=0.142$; Fig. 3b). Combined with the observed increase in individual connectivity (Fig. 3a), this pattern indicates that participants tended to consolidate into a large group, while some drifted away from the largest component. This consolidation was most pronounced in the individual assistance condition, which exhibited the highest overall connectivity.
Stance-based group composition, measured by stance assortativity (Fig. 2), can also be characterized by average stance differences within and between groups. We found that within-group stance distance remained stable or decreased slightly in the no-assistance and individual assistance conditions. In the relational assistance condition, within-group stance distance marginally increased relative to the no-assistance condition ($p=0.070$; Fig. 3c). In contrast, between-group stance distance increased significantly over rounds in the control condition ($p=0.017$), and this separation was further amplified under individual assistance ($p=0.004$; Fig. 3d). Relational assistance, however, did not differ significantly from the control in between-group stance distance ($p=0.350$). In summary, relational assistance increased ideological diversity within groups, whereas individual assistance widened separation between groups.
Figure 3: Structural dynamics with and without AI assistance. Changes from the initial round are shown for (a) number of conversation partners, (b) number of conversation groups, (c) within-group stance distance, and (d) between-group stance distance across conditions. Error bars represent mean $\pm$ s.e.m.
2.2 From AI Assistance to Conversational Receptiveness
AI assistance in this study operated solely at the level of individual message writing, without directly shaping or suggesting group formation. This implies that a mediating process likely links individual use of AI assistance to macro-level changes in social cohesion. We therefore examined participants’ AI suggestion use and group conversational dynamics as potential mediators.
In the AI-assisted conditions, participants continuously received message suggestions as the group chat unfolded and occasionally chose to send them. Overall, $45.1\%$ of messages in the individual condition and $37.0\%$ in the relational condition originated from GPT-4o. Most of these AI messages were sent without edits ($96.3\%$ of suggestions used). This AI use increased participants’ overall communication volume (Fig. 4a). When suggestions were available, participants sent more messages, an effect especially pronounced in the individual condition (up 37.9% in the individual condition; up $20.0\%$ in the relational condition; $p\leq 0.001$ for both). They, however, produced fewer self-written messages than in the no-assistance condition: the number of self-written messages decreased by $24.2\%$ in the individual assistance condition and by $24.5\%$ in the relational assistance condition ($p<0.001$ for both).
Figure 4: Relational assistance increased message receptiveness, while individual assistance sustained communication volume. Individual-level message counts (a) and sentiment (b), shown with and without AI-assisted messages. Group-level changes in message counts (c) and sentiment (d) across rounds. Error bars represent mean $\pm$ s.e.m.; $p$-values from regression analyses.
AI assistance also altered participants’ communication style (Fig. 4b). Relational assistance, which incorporated social awareness by design, generated more emotionally positive and receptive messages than individual assistance or self-written messages (Extended Data Fig. 1). As participants occasionally used these suggestions, the overall emotional valence of their messages increased; VADER scores, which quantify the sentiment polarity of text—with higher scores reflecting more positive emotional tone [35], rose from $0.113$ without assistance to $0.249$ with relational assistance ($p<0.001$). Notably, even participants’ self-written messages in the relational assistance condition were significantly more positive than those in the no-assistance condition ($p=0.006$). We confirmed that this increase in emotional positivity was correlated with other dimensions of conversational receptiveness [36, 37], including greater politeness (Extended Data Fig. 2) and lower toxicity (Extended Data Table 4).
Individual-level changes in messaging with AI assistance shaped group-level conversational dynamics (Extended Data Table 5). In the absence of AI assistance, groups had less communication as rounds progressed ($p=0.048$; Fig. 4c). When individual message suggestions were available, however, groups maintained communication volume, significantly reversing the decline ($p=0.008$). On the other hand, individual assistance did not affect group-level emotional valence ($p=0.206$; Fig. 4d). Relational assistance exhibited the opposite pattern: the volume of the message did not differ from the default ($p=0.664$), but baseline emotional valence was significantly higher ($p=0.004$) and remained stable across rounds.
Figure 5: Heavier AI users sent more messages. Dots represent individual participants’ AI suggestion use rate and their message counts per round under individual and relational assistance conditions ($N=182$ in the individual assistance condition; $N=188$ in the relational assistance condition). $r$ indicates Spearman’s rank correlation; both correlations are significant at $p<0.001$.
Regarding the individual–group effect, it is important to note that contributions to group discussions were not evenly distributed across participants. Unlike dyadic conversations driven by reciprocal turn-taking, group discussions are often shaped by a few members who contribute more frequently and, in doing so, set the conversational tone [38, 39]. We find that this heterogeneity was strongly associated with AI usage: participants who relied more heavily on AI message suggestions sent more messages to their groups. Figure 5 shows a strong positive correlation between AI suggestion use and message volume in the individual assistance condition (Spearman’s rank correlation $r=0.536$; $p<0.001$) and a moderate positive correlation in the relational assistance condition ($r=0.387$; $p<0.001$). This pattern likely reflects the reduced effort required to send AI-generated messages, enabling heavier users to contribute at a higher rate. As a result, the linguistic tone embedded in AI-generated suggestions (Extended Data Fig. 1) disproportionately influenced group conversations through these more AI-dependent participants.
2.3 From Conversational Receptiveness to Social Cohesion
After each discussion round, on average $32.6\%$ of participants left their groups to join another or create a new one. There were no significant differences in group-switching rates across conditions ($31.5\%$ in the control; $33.3\%$ in the individual condition; $33.1\%$ in the relational condition; one-way ANOVA, $p=0.553$). This suggests that the system-level differences observed in Figs. 2 and 3 reflect not the overall likelihood of switching groups, but by which groups they chose to leave and join.
Participants’ experiences during group discussions likely influenced whether they stayed in their current group or chose a new one, thereby linking micro-level interaction patterns to macro-level group composition Fig. 2. To examine this, we analyzed whether each participant’s message counts and sentiment predicted the stance distance between that participant and their group members in the next round, while controlling for their current stance distance (Fig. 6a). Linear mixed models were used for this analysis, with effects estimated separately for participants who stayed in the same group and those who switched groups (see Methods for details).
Figure 6: Effects of message sentiment and volume on ego-centric stance distance. (a) Example of changes in ego-centric stance distance across rounds. One member left and two joined (indicated by bold lines) from Round $t$ to Round $t+1$. Node color indicates each member’s stance (see Fig. 1). (b) Heatmap shows predicted between-round change in ego-centric stance distance for participants who remained in the same group, estimated from a linear mixed model (see Extended Data Table 6). The contour line marks the transition point ($\Delta=0$) between reduced and increased diversity. Axes represent average sentiment (VADER score) and number of messages sent. Subfigures show the density distributions of message sentiment and counts across AI assistance conditions.
Figure 6b presents the estimated results for participants who remained in the same group (see Extended Data Table 6 for model estimates). We found that emotional valence increased each participant’s stance distance when interacting with higher message counts ($p=0.031$). Given that participants rarely updated their stance during the conversation (6.40% of cases), this increase likely reflects the formation of more heterogeneous connections within the group rather than the emergence of polarization (Fig. 6a). In other words, when individuals sent more emotionally positive messages, their groups tended to attract a more ideologically diverse set of members. By contrast, for participants who switched groups, we observe no consistent relationship between prior communication and the composition of their subsequent group (Extended Data Table 6).
Finally, we examined the entire causal pathways using structural equation modeling [40] at the session level (Extended Data Table 7). The path diagram in Figure 7 reveals distinct mechanisms for individual and relational assistance. Individual assistance increased message volume, whereas relational assistance strongly increased message sentiment. Higher message volume alone was associated with greater stance assortativity, but its interaction with positive sentiment predicted lower assortativity in the subsequent round, reflecting the emergence of more heterogeneous group configurations. The effects of assistance on assortativity were fully mediated by conversational dynamics, indicating that structural changes emerge indirectly through interaction patterns rather than direct intervention.
Figure 7: Structural equation model of session-level effects of AI assistance on stance assortativity. Message volume, sentiment, and their interaction predict stance assortativity in the subsequent round, controlling for prior assortativity. The effects of individual and relational assistance are assessed relative to the no-assistance control condition. Solid arrows indicate standardized path coefficients, and dashed lines indicate covariances among mediators. Black lines indicate paths siginificant at $p<0.001$, and gray lines indicate paths that do not meet this threshold ($N=60$). Covariance estimates and indirect effects are provided in Extended Data Table 7.
3 Discussion
Our experiment demonstrates how LLM-powered message suggestions produce divergent effects on group composition dynamics in online discussions. Individual assistance, which tailored suggestions to users’ own stances, increased communication volume and group size but widened stance gaps across groups. Relational assistance, which incorporated awareness of other group members’ perspectives, fostered more emotionally positive and receptive exchanges, encouraging participants to form groups with greater opinion diversity. These contrasting pathways illustrate how subtle design choices in AI assistance can either entrench division or actively cultivate cohesion at scale.
Addressing social division and polarization is widely recognized as a pressing societal challenge [16, 1, 3, 29]. Prior work has tended to treat individual-level and system-level interventions separately [1, 41] with a limited understanding of how effects at one level propagate to the other. At the individual level, studies frequently focus on dyadic communication and examine how people might be guided toward consensus on controversial topics [42, 43, 27]. At the system level, interventions aim to increase opportunities for diverse populations to interact, such as by designing platforms that encourage cross-cutting exposure [32, 44]. However, these approaches remain disconnected: Individual-level studies lack mechanisms to evolve or even sustain opportunities for ongoing interaction [3], while system-level interventions often fall short because mere exposure is insufficient for social cohesion and can even backfire [32].
Our work complements these studies by explicitly addressing the dynamics between individual behavior and collective structure [6]. Rather than aiming for immediate consensus at the interpersonal level [29], we examine how the form of conversation can shape group communication norms and foster connections across differences [45]. Cohesion at the system level emerged through spontaneous interaction among individuals exercising free choice [46, 47]. Prior empirical work shows that certain linguistic patterns facilitate the formation of echo chambers online [48, 49]. Our findings suggest that targeted linguistic interventions can reverse this tendency, counteracting polarization dynamics and fostering more heterogeneous and cohesive group structures.
This study also highlights the risks and potential of AI assistance in shaping patterns of social division and cohesion. As AI technologies continue to advance, they have become embedded in everyday communication tools and are often personalized to individual users [21, 22]. While such personalization can improve usability, sycophantic AI assistance that echoes users’ views [26] may also amplify perceptual asymmetries, where “in-group virtues become out-group vices” [50].
This conventional personalization paradigm reflects methodological individualism in social science, which treats individuals as isolated units of analysis [51]. Our findings suggest complementing this perspective with a relational lens [52]. Individuals are not fixed entities but enact different roles and identities across social contexts [53, 54, 55]. If we view a “person” not only as an isolated individual but as someone embedded in interactional networks [56], AI personalization can be designed to help users express their opinions in ways that remain open and responsive to others—even those who disagree [57].
Human agency is another key consideration in the use of AI assistance. In our study, participants sent more messages when suggestions were available but produced fewer self-written ones, and they rarely edited the suggested content. These patterns suggest that participants delegated part of their agency to AI, primarily to reduce the effort of composing messages themselves [24]. As a result, relational assistance facilitated more receptive conversations and greater social cohesion, but these outcomes could occur without participants’ explicit intention. This dynamic resembles classic solutions to collective action problems, in which individual incentives help achieve public goods [58]. Our findings further suggest that humans may learn from AI by adopting more receptive language, as self-written messages became more emotionally positive when exposed to relational suggestions [27]. At the same time, however, the results raise concerns that human agency and interpersonal trust may erode as control over social interactions shifts toward automated systems [22, 59, 60, 61].
We acknowledge several limitations of this study that point to directions for future work. First, our experiment took place in a controlled online setting with structured discussions, whereas real-world conversations unfold over longer time horizons across multiple platforms [62]. Participants in our study rarely changed their opinion stance within a brief session, yet over extended interactions, established clusters may further polarize or, conversely, moderate through repeated exposure and trust-building [63, 64]. Second, future work can also advance the design space of conversational assistance itself. Although relational assistance effectively reduced stance-based assortativity, it was used less frequently than individual assistance, which led to higher conversational engagement. Exploring additional design dimensions—such as embodiment and controllability—may help balance social receptiveness with engagement [65, 66]. Finally, beyond technical refinement, evaluating these systems in politically salient contexts will be crucial for understanding how AI assistance operates amid the social and emotional complexities of real-world discourse [17, 15].
Taken together, this study demonstrates how subtle, AI-mediated changes in communication can scale from individual interactions to collective structures. By tracing the micro–macro loop, we show that message-level interventions can cascade into group-composition dynamics, either reinforcing divisions or fostering cohesion in opinion-diverse populations. As AI tools become increasingly ubiquitous, they offer an opportunity to redirect their influence toward the production of public goods. Reimagining personalization and autonomy as socially embedded processes may help resist the drift toward fragmentation and polarization that increasingly characterizes many societies today. Our findings suggest design levers for AI systems and communication policies that could facilitate more cohesive and resilient public discourse.
4 Methods
4.1 Ethics, Consent, and Preregistration
This study was preregistered [30] and approved by the Carnegie Mellon University Institutional Review Board. Participants provided informed consent prior to participation. Each participant received US\$6.50 for an approximately 30-minute task. No identifying information was collected.
4.2 Participants
We conducted the online experiment from July to August 2025 with 557 unique participants recruited from Prolific [31] across 60 sessions. Participants interacted anonymously via their web browser on a custom chat platform developed with Empirica [67]. Each individual could participate in only one session to avoid potential learning effects.
To increase the likelihood that participants shared context on the discussion topics, recruitment was restricted to U.S.-based participants. Demographic data, collected through a post-task survey response, are reported in Extended Data Table 8. Before participating, all individuals completed a human verification check and passed a comprehension test on the experimental procedure. The full task instructions are provided in the Supplementary Information.
4.3 Procedure
Upon entry, participants completed a short tutorial and a pre-task survey with seven items, each asking for agreement with a political statement on a 7-point Likert scale (Extended Data Table 1). Afterwards, they entered a virtual waiting room (up to 8 minutes) until at least twelve individuals were available, after which they were randomly assigned to one of the conversation groups (i.e., chatrooms).
Initial group size ranged from two to four participants. For example, in a session with 11 participants, they were assigned to three groups of size 4,4, and 3 (Fig. 1). Each session’s discussion topic was selected based on the greatest divergence (i.e., highest entropy) in participants’ pre-task survey responses. All groups within a session discussed the same topic, but consisted of different members. Participants could update their opinion stances (i.e., survey responses) during the session, and their stances were not visible to other participants.
Each session began with a 1-minute warm-up introduction, followed by four conversational rounds of 3 minutes each. In each round, participants engaged in free-form discussion on the assigned topic. Participants were anonymous to one another and identifiable only by self-selected ad-hoc identifiers. In the AI assistance conditions, participants received real-time message suggestions displayed above the text-entry field as the group conversation unfolded (see SI). They could send these suggestions directly with a click, copy them into the text box and edit before sending, or compose their own messages while suggestions were displayed.
After each round, participants had 1 minute to switch groups before the next round, enabling dynamic reconfiguration of their group membership. During this period, they could remain in the current group, switch to another group, or create a new group, while viewing the conversation history and current membership of other groups. To support informed decision-making, they were guided through this process individually during the tutorial. Occasionally, only one participant remained in a group; in such cases, they continued the discussion with a chatbot to ensure that everyone had an active partner. The chatbot generated responses using GPT-4o, independently of the message-suggestion system.
At the end of the session, participants completed a post-task survey about their experience and basic demographics. They then received a base payment of \$5.00 for completing the study, with an additional \$1.50 performance bonus contingent on active engagement in the conversation.
4.4 Message Suggestions
In the AI assistance conditions, we used OpenAI’s GPT-4o model to generate real-time message suggestions during the conversation. The prompts were designed to increase the likelihood that suggestions would be adopted by users by incorporating the preceding chat history, specifying a conversation strategy, and producing concise, relevant messages (see SI for full prompts).
In the individual assistance condition, the model was further prompted to align suggestions with each participant’s stance on the discussion topic. Specifically, we instructed the model—such as “The content should align with their opinion rating, reinforcing their stance (even if it is divisive). Emphasize their perspective firmly rather than seeking compromise.”—to generate messages personalized to each individual.
In contrast, in the relational assistance condition, the model was prompted to consider the stances of other group members in addition to the focal participant’s stance. This design draws on theories of communication accommodation, and relational framing, which argue that conversational style shifts depending on the relationships among participants [68, 69, 70]. We operationalized these theoretical ideas into relational assistance by using prompt instructions such as “The content should align with (user’s ID) ’s viewpoint while acknowledging others’ points, finding common ground, or using a cooperative tone.” Unlike individual assistance, the generated suggestions in this condition varied dynamically with the group composition the participant belonged.
All participants in a given session were assigned to the same assistance condition. Extended Data Table 2 illustrates differences in suggested messages across conditions resulting from this prompt tuning.
4.5 Analyses
4.5.1 Group Composition
To analyze group composition dynamics, we constructed a bipartite network representation of each session round and projected it onto a participant–participant network, where an edge existed if two participants were in the same group (Fig. 1) [33]. Each participant‘s opinion stance was quantified on a scale from –3 (strongly disagree) to +3 (strongly agree), based on their 7-point Likert scale responses to the discussion topic. From this structure, we derived four measures of social organization: stance assortativity, the number of conversation partners, the number of conversation groups, within-group stance distance, and between-group stance distance.
Stance assortativity was computed to assess the extent to which conversational ties (i.e., co-membership in a group) connected participants with similar versus dissimilar stances. Stance assortativity values range from –1 (complete disassortativity) to +1 (complete homophily), with 0 indicating random mixing relative to the distribution of stances [34].
Number of partners was defined as the number of unique conversation partners a participant had within their group in a given round, regardless of stance. This measure captures the size of each participant’s ego-centric network and reflects how group reconfigurations expanded or reduced opportunities for interaction. Number of groups, in contrast, was defined as the total number of conversation groups per session in each round. This measure indicates the degree of structural divergence or convergence at the system level.
Within-group stance distance was measured as the mean absolute difference in stance scores among members of the same group, capturing the level of opinion diversity within groups. In contrast, Between-group stance distance was measured as the mean absolute difference in average stance scores between groups, capturing separation at the system level.
Because random assignment produced variation in initial group size and stance distribution, these measures varied across sessions and conditions by chance. To enable clearer comparisons, we examined structural measures from Round 2, subtracting the first-round baseline.
4.5.2 Conversation Dynamics
To analyze conversational patterns within groups, we focused on participants’ message-level behaviors. We measured individual-level messaging activity and the use of AI-generated suggestions.
Message volume was defined as the total number of messages sent by each participant in a round, capturing their level of engagement in group discussion. As an indicator of conversational style, we estimated message sentiment using the VADER sentiment analysis toolkit [35], which produces a compound score representing the overall polarity of text. Higher scores indicate a more positive and receptive emotional tone. VADER is well-suited for the context of this study, which involves multiperson conversations and online discourse. Sentiment scores were averaged across all messages a participant sent within each round. To complement this analysis, we also applied additional toolkits [36, 37] to assess other dimensions of conversational receptiveness, including politeness and toxicity.
In the AI assistance conditions, we also calculated these measures for self-written messages separately. We identified a message as AI-originated if (1) it was sent directly using the suggestion accept button, or (2) it was copied from the suggestion into the text box and then sent, with or without edits. We then analyzed the volume and sentiment of the messages for the remaining self-written messages.
We also aggregated these measures at the group level across rounds and conditions. For each group, we counted the total number of messages and averaged the VADER scores of the entire group conversation to estimate group-level sentiment (Fig. 4cd).
4.5.3 Statistical Modeling from Conversation to Cohesion
To examine how conversational behaviors predicted subsequent ego-centric stance distance within groups, we estimated linear mixed models at the participant–round level. Ego-centric stance distance was defined as the mean absolute difference in stance between a focal participant and their group members. We conducted separate analyses for participants who remained in the same group across consecutive rounds and those who switched groups, thereby isolating the effects of ongoing communication from those of group reconfiguration.
The dependent variable was the participant’s ego-centric stance distance in the next round. The model included the following predictors:
  • ego-centric stance distance in the current round (baseline control),
  • average message sentiment (VADER score),
  • message volume (number of messages sent), and
  • the interaction between sentiment and message volume.
Random intercepts were specified for participants nested within groups to account for repeated observations and group-level dependencies. Formally, the model was expressed as:
$\text{Stance Distance}_{i,t+1}$ $=\beta_{0}+\beta_{1}\,\text{Stance Distance}_{i,t}+\beta_{2}\,\text{Message Sentiment}_{i,t}$
$\quad+\beta_{3}\,\text{Message Volume}_{i,t}+\beta_{4}\,(\text{Message Sentiment}_{i,t}\times\text{Volume}_{i,t})$
$\quad+u_{\text{participant}}+u_{\text{group}}+\epsilon$
where $u_{\text{participant}}$ and $u_{\text{group}}$ represent random intercepts for participants and groups, respectively. This model tested whether more frequent and more emotionally positive communication predicted changes in ego-centric stance distance, above and beyond baseline differences. Extended Data Table 5 reports the estimated coefficients. Figure 6 visualizes the predicted stance distance based on these estimated values for participants who remained in their groups.
Finally, to examine the entire causal pathways, we conducted structural equation modeling (SEM) at the session level [71]. This analysis allowed us to capture how AI assistance conditions influenced stance assortativity through their effects on conversational mediators. Specifically, we modeled stance assortativity in Round $T+1$ as a function of:
  • stance assortativity in Round $T$ (baseline control),
  • average message sentiment,
  • message volume,
  • the interaction between sentiment and volume, and
  • individual and relational assistance (relative to the no-assistance control condition).
Message volume and sentiment were computed using normalized (but not centered) values to preserve the interpretability of zero as a meaningful baseline. The model included both direct paths from the experimental conditions to stance assortativity and indirect paths mediated by conversational dynamics. Covariances among mediators (sentiment, volume, and their interaction) were freely estimated to account for their interdependence. Indirect effects were computed to assess how AI assistance influenced stance assortativity through changes in conversational tone and activity. Figure 7 presents the resulting path diagram, and Extended Data Table 7 reports the full parameter estimates.
[…]
Acknowledgements We thank Y. Li, A. Nairo, S. Rathje, N. A. Christakis, and R. E. Kraut for their insightful feedback on the manuscript. This work was supported by National Science Foundation (grant no 2237095) and the NOMIS foundation.
Declarations
[…]
Funding National Science Foundation (#2237095) and the NOMIS foundation.
[…]
Competing interests The authors declare no competing interests.
[…]
Data availability The data generated and analyzed will be available upon publication.
[…]
Contributions F.H.: Methodology, Software, Data Collection, Analysis, Writing. E.L.C.: Conceptualization, Methodology, Software, Writing. H.S.: Conceptualization, Methodology, Analysis, Writing, Funding Acquisition. All authors read and approved the final manuscript.

### v2

@@
-From Social Division to Cohesion with AI Message Suggestions in Online Chat Groups
+AI-Mediated Communication Reshapes Social Structure in Opinion-Diverse Groups
 Abstract
-Social cohesion is difficult to sustain in societies marked by opinion diversity, particularly in online communication.
-As large language model (LLM)-driven messaging assistance becomes increasingly embedded in these contexts, it raises critical questions about its societal impact.
-We present an online experiment with 557 participants who engaged in multi-round discussions on politically controversial topics while freely reconfiguring their discussion groups.
-In some conditions, participants received real-time message suggestions generated by an LLM, either personalized to the individual or adapted to their group context.
-We find that subtle shifts in linguistic style during communication, mediated by AI assistance, can scale up to reshape collective structures.
-While individual-focused assistance leads users to segregate into like-minded groups, relational assistance that incorporates group members’ stances enhances cohesion through more receptive exchanges.
-These findings demonstrate that AI-mediated communication can support social cohesion in diverse groups, but outcomes critically depend on how personalization is designed.
+Group segregation or cohesion can emerge from micro-level communication, and AI-assisted messaging may shape this process.
+Here, we report a preregistered online experiment (N = 557 across 60 sessions) in which participants discussed controversial political topics over multiple rounds and could freely change groups.
+Some participants received real-time message suggestions from a large language model (LLM), either personalized to their stance (“individual assistance”) or incorporating their group members’ perspectives (“relational assistance”).
+We find that small variations in AI-mediated communication cascade into macro-level differences in group composition.
+Participants with individual assistance send more messages and show greater stance-based clustering, whereas those with relational assistance use more receptive language and form more heterogeneous ties.
+Hybrid expressive processes—jointly produced by humans and AI—can reshape collective organization.
+The patterns of structural division and cohesion depend on how AI incorporates users’ interaction context.
 keywords:
 social cohesion, opinion diversity, online communication, large language models, human–AI interaction
-1 Introduction
-Social cohesion remains a persistent challenge in societies characterized by opinion diversity [1, 2, 3].
-Conversations across differences are fraught: people struggle to engage constructively with those who hold opposing views, and stereotypes bias perceptions of outgroup members [4, 5].
-The challenges individuals face in conversation are reinforced by structural dynamics, as homophily drives clustering with similar others and limits cross-cutting ties at scale [6, 7, 8].
-Online platforms intensify this micro–macro loop by lowering inhibitions in communication, accelerating interpersonal clustering, and amplifying these effects at scale [9, 10, 11, 12, 13].
-The consequences—from political polarization to breakdowns in global coordination—underscore the urgency of understanding how opinion diversity is managed within online environments [14, 15, 16, 17, 18].
-Recent advances in artificial intelligence (AI) introduce a new layer of influence to this challenge [19, 20].
-Large language models (LLMs) are now widely integrated into everyday online communication through messaging assistance tools [21, 22, 23].
-While such assistance can reduce the burden of writing [24], conventional personalization features risk amplifying stance differences among people with diverse views [25, 26].
-Conversely, some designs for AI-mediated communication have been proposed to foster interpersonal trust and build consensus across lines of disagreement [27, 1, 22, 28, 29].
-Yet, it remains unclear how these micro-level interventions can cascade upward to shape macro-level changes in human group composition.
-How can AI message suggestions foster social cohesion rather than exacerbate division in diverse populations?
-Here, we address this question through a preregistered experiment conducted on a customized online chat platform (see Methods for details) [30].
+Understanding how micro-level communication patterns accumulate into macro-level group segregation or cohesion is a central question in social and behavioral science [1, 2, 3].
+Conversations across differences are often asymmetric: people find it difficult to engage constructively with those who hold opposing views [4, 5], and stereotypes bias perceptions of outgroup members [6].
+These interpersonal tendencies interact with structural forces—particularly homophily and clustering—to produce emergent segregation, even without explicit hostility [2, 7, 8].
+Online platforms can intensify these dynamics through lowered inhibitions [9], emotion-amplified diffusion [10], and algorithmic or behavioral clustering processes [11, 12, 13].
+While the forces that produce social division are well theorized and empirically documented, far less is known about the micro-level conversational mechanisms that can instead generate cohesion in ideollogically diverse groups [14, 15, 16].
+Recent advances in artificial intelligence (AI) introduce a new layer of influence to these micro-macro dynamics [17].
+Large language models (LLMs) are increasingly integrated into everyday communication through messaging assistance tools [18, 19].
+These tools effectively create hybrid human–AI expressive processes, in which algorithmic suggestions subtly shape how people articulate their views [20, 21, 22, 23].
+While such assistance can reduce the burden of writing [24], conventional personalization features may unintentionally reinforce users’ ideological viewpoints and deepen intergroup divisions [25, 26].
+In contrast, recent work shows that assistance incorporating the perspectives of interaction partners can improve cross-group receptiveness [27, 28, 29, 30].
+Yet it remains unclear whether—and how—such micro-level shifts in AI-mediated expression scale up to shape macro-level patterns of group composition in hybrid communication environments.
+To examine these dynamics, we conducted a preregistered experiment on an online chat platform (see Methods for details) [31].
 Participants ($N=557$), recruited from Prolific [32], were randomly assigned to sessions of 6 to 15 people ($M=9.42$, $s.d.=2.57$), subdivided into groups of 2 to 4 (Fig.
 1).
-Each group discussed a politically controversial topic selected to maximize opinion diversity for that session, based on participants’ pre-task survey responses (see Extended Data Table 1 for survey items) [27].
-Participants were able to update their survey stance during the session, though this occurred rarely ($6.40\%$ of cases).
-Participants were instructed to engage in group conversations, but were not asked to reach consensus, reflecting the open-ended and non-teleological nature of informal online discussions [32, 14].
-After each three-minute discussion, participants decided whether to remain in their current group, join another, or create a new one.
-If only one participant remained in a group, they continued the discussion with a chatbot to ensure continuous engagement.
-This procedure was repeated for four rounds, followed by a final switching decision that determined the Round 5 group composition used for structural analysis.
-Within this setup, we manipulated the presence and type of AI assistance in message writing through three experimental conditions [27].
-In the no-assistance condition, participants composed all messages themselves and did not receive any AI-generated suggestions.
-In the other two conditions, participants received real-time message suggestions generated by OpenAI’s GPT-4o model as the group chat unfolded.
-Although suggestions were tailored to each user based on their survey responses and subsequent updates in both conditions, the personalization context differed between them (see Methods; full prompts appear in SI.).
-In the individual assistance condition, the model was prompted to help participants articulate their stance on the conversation topic, aligning suggestions with each individual’s viewpoint.
-In contrast, in the relational assistance condition, the modelincorporated not only the participant’s stance but also those of their group members, with the goal of fostering mutual understanding.
-As a result, individual-focused suggestions primarily supported self-expression, whereas relationally aware suggestions more often conveyed receptive language and positive sentiment, even in the presence of disagreement (Extended Data Table 2), as confirmed by quantitative analyses (Extended Data Fig.
+Groups discussed a politically controversial topic selected to maximize opinion diversity within each session, based on pre-task survey responses (Extended Data Table 1) [27].
+Participants could update their stance during the session, though this occurred rarely ($6.40\%$ of cases).
+Participants were instructed to engage in open-ended discussions, without any requirement to reach consensus, mirroring the non-teleological nature of informal online conversations [33, 13].
+After each three-minute discussion, participants chose whether to remain in their current group, join another, or create a new one.
+If a group had only one participant, they continued the discussion with a chatbot to ensure continuous engagement.
+This process was repeated for four rounds, followed by a final switching decision that determined the final group configuration used for structural analysis.
+Within this setup, we manipulated the presence and construction of AI assistance through three conditions [27].
+In the no-assistance condition, participants composed all messages themselves.
+In the other two conditions, participants received real-time message suggestions generated by OpenAI’s GPT-4o model, which they could send directly with one click, edit, or ignore.
+Suggestions were personalized to each participant based on their survey-reported stance and the unfolding conversation, but the nature of this personalization differed across the two assisted conditions.
+In the individual assistance condition, the model was prompted to help participants articulate their own stance on the discussion topic, supporting self-expression.
+In the relational assistance condition, the model was prompted to incorporate both the participant’s stance and those of their group members, with the emphasis on fostering mutual understanding.
+Individual and relational prompts were matched in length and complexity and differed only in instructional framing (see Methods; full prompts appear in SI).
+As a consequence, individual-focused suggestions emphasized self-expression, whereas relationally framed suggestions adapted to group composition and conveyed more positive sentiment, even in the presence of disagreement (Extended Data Table 2), a pattern confirmed by quantitative analyses (Extended Data Fig.
 1).
-These suggestions could be sent with a single click, edited, or ignored in favor of self-written messages.
+This variation enabled a causal test of how distinct expressive framings influence group-level conversation and structural dynamics.
 Following the preregistered protocol [31], we excluded sessions with fewer than six participants or with more than 25% dropout.
-The final dataset includes 20 sessions per condition (60 sessions total), comprising 557 participants overall (187 in the no assistance condition, 182 in the individual assistance condition, and 188 in the relational assistance condition).
-Each participant completed only one session.
-Per-session participant counts, dropout rates, and topic assignments did not differ significantly across conditions (Chi-square tests for participant counts: $p=0.946$; dropout rates: $p=0.538$; Fisher’s exact test for topic distributions: $p=0.533$).
-We evaluate the effects of AI assistance on group composition and social cohesion, measured through stance assortativity—a session-level indicator of how strongly group composition reflects similarity in participants’ opinions.
-Individuals are treated as connected if they belong to the same group [33], and assortativity is computed from their stances on the discussion topic (3-3 = totally disagree to +3+3 = totally agree) [34].
-Higher assortativity values indicate stronger clustering of like-minded individuals (i.e., greater social division and lower cohesion), whereas lower values indicate more cross-cutting ties (i.e., lower division and greater cohesion) (Fig.
+The final dataset includes 20 sessions per condition (60 sessions total), comprising 557 unique participants overall (187 in the no assistance condition, 182 in the individual assistance condition, and 188 in the relational assistance condition).
+Per-session participant counts, dropout rates, and topic assignments did not differ significantly across conditions (participant counts: $p=0.946$; dropout: $p=0.538$; topics: $p=0.533$).
+We evaluate effects on group composition using stance assortativity, a structural measure of how strongly group membership reflects similarity in opinions (Fig.
 1).
+Individuals are defined as connected if they belong to the same group [34], and assortativity is computed from their stance values (3-3 = totally disagree to +3+3 = totally agree) [35].
+Higher assortativity indicates stronger clustering of like-minded individuals, whereas lower assortativity indicates more cross-cutting ties and greater cohesion.
 Because random assignment produced variation in initial group composition, we analyze changes relative to the first-round baseline.
-To examine the mechanisms underlying these structural shifts, we also analyze conversational dynamics, including message volume and sentiment, with the latter measured using VADER sentiment analysis [35].
-Figure 1: Illustration of group-composition dynamics.
-Each session begins with 2 to 4 groups, each consisting of 2 to 4 members, shown as gray clusters.
-Stance assortativity is calculated at the session level from participants’ stances on the discussion topic, represented by node colors.
-As participants switch groups across rounds, group composition evolves, either increasing social division (positive assortativity) or fostering cohesion (negative assortativity).
-2 Results
-2.1 Effects of AI Assistance on Social Cohesion
+To probe the mechanisms underlying these structural shifts, we also examine conversational dynamics, including message volume and sentiment (measured using VADER [36]).
+Figure 1: AI-mediated communication and group reconfiguration.
+This schematic illustrates the experimental setup and the conceptual process by which group composition can reorganize during the study.
+Nodes represent participants (with colors indicating their stance), and edges denote co-membership in discussion groups.
+In the AI-assisted conditions, message suggestions generated by a LLM are tailored to each participant’s stance and conversational context, creating hybrid human–AI expressive processes.
+Stance assortativity is computed at the session level and reflects how strongly group membership clusters like-minded individuals.
+As participants switch groups, group composition can reorganize in ways that increase or decrease assortativity, corresponding to greater social division or cohesion.
+Results
+Effects of AI Assistance on Group Reconfiguration
 Figure 2: Relational AI assistance reduces stance assortativity.
@@
 Error bars represent mean $\pm$ s.e.m.
-We found that stance assortativity remained stable or increased slightly across rounds in the no-assistance condition, but decreased substantially in the relational assistance condition (Fig.
+We found that stance assortativity remained stable or increased slightly across rounds in the no-assistance condition, but decreased substantially under relational assistance (Fig.
 2).
-After four rounds, assortativity decreased by an average of 0.135 points on the –1 to 1 scale under relational assistance.
-This reduction was statistically significant relative to the no-assistance condition ($p=0.004$; regression analysis; Extended Data Table 3).
+After four rounds, assortativity decreased by an average of 0.135 points (on the –1 to 1 scale) in the relational assistance condition, a statistically significant reduction relative to the no-assistance control ($p=0.004$; regression analysis; Extended Data Table 3).
 In contrast, changes in assortativity under individual assistance did not differ significantly from the control ($p=0.608$).
-Relational assistance promoted more cross-cutting group composition, while individual assistance did not alter baseline patterns of clustering.
-To further understand the effects of AI assistance on structural dynamics, we examined the number of conversation partners, the number of conversation groups, within-group stance distance, and between-group stance distance across interaction rounds.
-Participants, on average, increased their conversation partners without AI assistance ($p<0.001$ in the no-assistance condition; Fig.
+Thus, relational assistance promoted more heterogeneous, cross-cutting ties, while individual assistance did not alter baseline assortativity.
+To further characterize how AI assistance shaped structural dynamics, we examined changes in the number of conversation partners and the number of conversation groups across rounds.
+Participants, on average, increased their number of conversation partners even without assistance ($p<0.001$ in the no-assistance condition; Fig.
 3a).
-This trend was observed across all conditions but it was amplified under individual assistance: the increase in conversation partners was significantly larger than in the control ($p=0.028$; Extended Data Table 3).
-Relational assistance did not differ significantly from the control ($p=0.712$).
-The number of groups remained stable across rounds in the control condition ($p=0.280$), and neither AI condition altered this pattern (individual: $p=0.317$; relational: $p=0.142$; Fig.
+This increase was significantly larger under individual assistance than in the non-assistance control ($p=0.028$; Extended Data Table 3), whereas relational assistance did not differ from the control ($p=0.712$).
+The number of groups remained stable across rounds in the control condition ($p=0.280$), and neither form of assistance significantly altered this pattern (individual: $p=0.317$; relational: $p=0.142$; Fig.
 3b).
-Combined with the observed increase in individual connectivity (Fig.
-3a), this pattern indicates that participants tended to consolidate into a large group, while some drifted away from the largest component.
-This consolidation was most pronounced in the individual assistance condition, which exhibited the highest overall connectivity.
-Stance-based group composition, measured by stance assortativity (Fig.
-2), can also be characterized by average stance differences within and between groups.
-We found that within-group stance distance remained stable or decreased slightly in the no-assistance and individual assistance conditions.
-In the relational assistance condition, within-group stance distance marginally increased relative to the no-assistance condition ($p=0.070$; Fig.
+Given that individual connectivity increased while the number of groups remained constant, participants tended to consolidate into a single large group while some drifted away from the largest component.
+This consolidation was most pronounced under individual assistance, which exhibited the highest overall connectivity.
+We further examined structural diversity by complementing stance assortativity with group-based measures of average stance differences within and between groups.
+Within-group stance distance remained stable or decreased slightly in the no-assistance and individual assistance conditions.
+Under relational assistance, within-group stance distance marginally increased relative to the no-assistance condition ($p=0.070$; Fig.
 3c).
-In contrast, between-group stance distance increased significantly over rounds in the control condition ($p=0.017$), and this separation was further amplified under individual assistance ($p=0.004$; Fig.
+Between-group stance distance increased significantly over rounds in the control condition ($p=0.017$), and this separation was further amplified under individual assistance ($p=0.004$; Fig.
 3d).
-Relational assistance, however, did not differ significantly from the control in between-group stance distance ($p=0.350$).
-In summary, relational assistance increased ideological diversity within groups, whereas individual assistance widened separation between groups.
+Relational assistance did not differ significantly from the control in between-group stance distance ($p=0.350$).
+Taken together, individual assistance produced groups that became more internally uniform and more externally separated, whereas relational assistance did not increase between-group separation and instead diversified group memberships.
 Figure 3: Structural dynamics with and without AI assistance.
@@
 Error bars represent mean $\pm$ s.e.m.
-2.2 From AI Assistance to Conversational Receptiveness
-AI assistance in this study operated solely at the level of individual message writing, without directly shaping or suggesting group formation.
-This implies that a mediating process likely links individual use of AI assistance to macro-level changes in social cohesion.
-We therefore examined participants’ AI suggestion use and group conversational dynamics as potential mediators.
-In the AI-assisted conditions, participants continuously received message suggestions as the group chat unfolded and occasionally chose to send them.
+From AI Assistance to Conversational Dynamics
+AI assistance in this study operated only at the level of individual message composition and never directly involved group reconfiguration.
+Accordingly, any macro-level structural differences must emerge through processes unfolding within the conversations themselves.
+To identify these processes, we examined participants’ use of AI suggestions and the resulting conversational dynamics as potential mediators linking expressive framing to group-level outcomes.
+In the AI-assisted conditions, participants continuously received message suggestions as the conversation unfolded and occasionally chose to send them.
 Overall, $45.1\%$ of messages in the individual condition and $37.0\%$ in the relational condition originated from GPT-4o.
-Most of these AI messages were sent without edits ($96.3\%$ of suggestions used).
-This AI use increased participants’ overall communication volume (Fig.
-4a).
-When suggestions were available, participants sent more messages, an effect especially pronounced in the individual condition (up 37.9% in the individual condition; up $20.0\%$ in the relational condition; $p\leq 0.001$ for both).
-They, however, produced fewer self-written messages than in the no-assistance condition: the number of self-written messages decreased by $24.2\%$ in the individual assistance condition and by $24.5\%$ in the relational assistance condition ($p<0.001$ for both).
+Moreover, 96.3% of these suggestions were sent without modification, indicating a high degree of reliance on and integration of AI-generated language within the hybrid human–AI expressive process.
+When suggestions were available, participants sent more total messages (Fig.
+4a)—an effect especially pronounced in the individual condition (up to a 37.9% increase in total messages; $p<0.001$) and to a lesser extent in the relational condition (up to a 20.0% increase; $p=0.001$).
+On the other hand, participants produced fewer self-written messages under AI assistance.
+Self-written message volume decreased by 24.2% in the individual condition and by 24.5% in the relational condition ($p<0.001$ for both).
 Figure 4: Relational assistance increased message receptiveness, while individual assistance sustained communication volume.
@@
 4b).
-Relational assistance, which incorporated social awareness by design, generated more emotionally positive and receptive messages than individual assistance or self-written messages (Extended Data Fig.
+Relational assistance, which incorporated social awareness by design, generated more emotionally positive and receptive suggestions than individual assistance or self-written messages (Extended Data Fig.
 1).
-As participants occasionally used these suggestions, the overall emotional valence of their messages increased; VADER scores, which quantify the sentiment polarity of text—with higher scores reflecting more positive emotional tone [35], rose from $0.113$ without assistance to $0.249$ with relational assistance ($p<0.001$).
+Participants who received these suggestions sent messages with significantly more positive emotional tone.
+VADER scores, which quantify the sentiment polarity of text, rose from $0.113$ to $0.249$ under relational assistance ($p<0.001$) (higher scores reflecting more positive emotional tone [36]).
 Notably, even participants’ self-written messages in the relational assistance condition were significantly more positive than those in the no-assistance condition ($p=0.006$).
-We confirmed that this increase in emotional positivity was correlated with other dimensions of conversational receptiveness [36, 37], including greater politeness (Extended Data Fig.
-2) and lower toxicity (Extended Data Table 4).
-Individual-level changes in messaging with AI assistance shaped group-level conversational dynamics (Extended Data Table 5).
-In the absence of AI assistance, groups had less communication as rounds progressed ($p=0.048$; Fig.
+We confirmed that this increase in emotional positivity was correlated with other dimensions of conversational receptiveness [37, 38], including greater politeness and lower toxicity (Extended Data Fig.
+2).
+Individual-level changes in messaging with AI assistance scaled up to shape group-level conversational dynamics (see Extended Data Table 4 for illustrative discussion excerpts).
+In the absence of AI assistance, groups communicated less as rounds progressed ($p=0.048$; Fig.
 4c).
-When individual message suggestions were available, however, groups maintained communication volume, significantly reversing the decline ($p=0.008$).
-On the other hand, individual assistance did not affect group-level emotional valence ($p=0.206$; Fig.
-4d).
-Relational assistance exhibited the opposite pattern: the volume of the message did not differ from the default ($p=0.664$), but baseline emotional valence was significantly higher ($p=0.004$) and remained stable across rounds.
+Under individual assistance, groups maintained conversation volume and significantly reversed this downward trend ($p=0.008$), while conversational sentiment remained unchanged ($p=0.206$).
+Relational assistance, in contrast, did not alter conversation volume trajectories ($p=0.664$), but produced significantly more positive baseline sentiment ($p=0.004$), which remained stable across rounds (Fig.
+4d; see Extended Data Table 5 for full statistical results).
 Figure 5: Heavier AI users sent more messages.
@@
 $r$ indicates Spearman’s rank correlation; both correlations are significant at $p<0.001$.
-Regarding the individual–group effect, it is important to note that contributions to group discussions were not evenly distributed across participants.
-Unlike dyadic conversations driven by reciprocal turn-taking, group discussions are often shaped by a few members who contribute more frequently and, in doing so, set the conversational tone [38, 39].
-We find that this heterogeneity was strongly associated with AI usage: participants who relied more heavily on AI message suggestions sent more messages to their groups.
-Figure 5 shows a strong positive correlation between AI suggestion use and message volume in the individual assistance condition (Spearman’s rank correlation $r=0.536$; $p<0.001$) and a moderate positive correlation in the relational assistance condition ($r=0.387$; $p<0.001$).
-This pattern likely reflects the reduced effort required to send AI-generated messages, enabling heavier users to contribute at a higher rate.
-As a result, the linguistic tone embedded in AI-generated suggestions (Extended Data Fig.
-1) disproportionately influenced group conversations through these more AI-dependent participants.
-2.3 From Conversational Receptiveness to Social Cohesion
-After each discussion round, on average $32.6\%$ of participants left their groups to join another or create a new one.
-There were no significant differences in group-switching rates across conditions ($31.5\%$ in the control; $33.3\%$ in the individual condition; $33.1\%$ in the relational condition; one-way ANOVA, $p=0.553$).
-This suggests that the system-level differences observed in Figs.
-2 and 3 reflect not the overall likelihood of switching groups, but by which groups they chose to leave and join.
-Participants’ experiences during group discussions likely influenced whether they stayed in their current group or chose a new one, thereby linking micro-level interaction patterns to macro-level group composition Fig.
-2.
-To examine this, we analyzed whether each participant’s message counts and sentiment predicted the stance distance between that participant and their group members in the next round, while controlling for their current stance distance (Fig.
+Regarding the individual–group link, contributions to group discussions were not evenly distributed across participants.
+Unlike dyadic conversations driven by reciprocal turn-taking, group discussions are often dominated by members who contribute more frequently and, in doing so, set the conversational tone and norms [39, 40].
+Figure 5 shows that this heterogeneity was strongly associated with AI usage: participants who relied more heavily on AI suggestions sent more messages to their groups (Spearman’s $r=0.536$ under individual assistance; $r=0.387$ under relational assistance; $p<0.001$ for both).
+This pattern likely reflects that the reduced effort required to send AI-generated messages enables heavier users to contribute at a higher rate.
+As a consequence, uneven reliance on AI created asymmetric influence within group discourse, amplifying the expressive impact of the most AI-dependent participants.
+From Conversational Dynamics to Group Reconfiguration
+After each discussion round, an average of $32.6\%$ of participants left their groups to join another group or create a new one.
+Group-switching rates did not differ significantly across conditions ($31.5\%$ in the no-assistance condtion; $33.3\%$ in the individual assistance condition; $33.1\%$ in the relational assistance condition; one-way ANOVA, $p=0.553$).
+This indicates that the system-level differences observed in Figs.
+2 and 3 were not driven by whether participants switched groups, but rather by where they moved.
+And participants’ conversational experiences likely influenced their subsequent group-switching decisions.
+To examine this linkage, we analyzed whether a participant’s message volume and sentiment in a given round predicted changes in their stance distance from their group members in the next round, controlling for their current stance distance (Fig.
 6a).
-Linear mixed models were used for this analysis, with effects estimated separately for participants who stayed in the same group and those who switched groups (see Methods for details).
+Figure 6b shows that more positive message sentiment increased next-round stance distance when paired with higher message volume ($p=0.031$; see Extended Data Table 6 for full model results).
+Because participants rarely updated their own stance relative to the pre-task survey (6.40% of cases), these changes primarily reflect shifts in group composition rather than ideological movement by individuals.
+In effect, participants who expressed more positive sentiment tended to end up in groups that became more ideologically diverse in the subsequent round.
+This provides a micro-level mechanism consistent with the reduction in stance assortativity observed under relational assistance (Fig.
+2).
 Figure 6: Effects of message sentiment and volume on ego-centric stance distance.
 (a) Example of changes in ego-centric stance distance across rounds.
-One member left and two joined (indicated by bold lines) from Round $t$ to Round $t+1$.
+From Round $t$ to Round t+1t{+}1, one member left and two members joined the group (bold edges).
 Node color indicates each member’s stance (see Fig.
 1).
-(b) Heatmap shows predicted between-round change in ego-centric stance distance for participants who remained in the same group, estimated from a linear mixed model (see Extended Data Table 6).
-The contour line marks the transition point ($\Delta=0$) between reduced and increased diversity.
+(b) Predicted between-round change in ego-centric stance distance for participants who remained in the same group, estimated using a linear mixed model (Extended Data Table 6).
+The contour line marks the transition point ($\Delta=0$) between reduced and increased stance distance.
 Axes represent average sentiment (VADER score) and number of messages sent.
-Subfigures show the density distributions of message sentiment and counts across AI assistance conditions.
-Figure 6b presents the estimated results for participants who remained in the same group (see Extended Data Table 6 for model estimates).
-We found that emotional valence increased each participant’s stance distance when interacting with higher message counts ($p=0.031$).
-Given that participants rarely updated their stance during the conversation (6.40% of cases), this increase likely reflects the formation of more heterogeneous connections within the group rather than the emergence of polarization (Fig.
-6a).
-In other words, when individuals sent more emotionally positive messages, their groups tended to attract a more ideologically diverse set of members.
-By contrast, for participants who switched groups, we observe no consistent relationship between prior communication and the composition of their subsequent group (Extended Data Table 6).
-Finally, we examined the entire causal pathways using structural equation modeling [40] at the session level (Extended Data Table 7).
+Marginal density plots show the observed distributions of message sentiment and message counts across the three AI assistance conditions.
+Finally, we examined the full causal pathways using structural equation modeling [41] at the session level (Extended Data Table 7).
 The path diagram in Figure 7 reveals distinct mechanisms for individual and relational assistance.
-Individual assistance increased message volume, whereas relational assistance strongly increased message sentiment.
+Individual assistance mainly increased message volume, whereas relational assistance increased message sentiment.
 Higher message volume alone was associated with greater stance assortativity, but its interaction with positive sentiment predicted lower assortativity in the subsequent round, reflecting the emergence of more heterogeneous group configurations.
-The effects of assistance on assortativity were fully mediated by conversational dynamics, indicating that structural changes emerge indirectly through interaction patterns rather than direct intervention.
+The effects of assistance on assortativity were fully mediated by conversational dynamics (i.e., no significant direct effects), indicating that structural changes arose indirectly through interaction patterns rather than direct intervention.
+Consistent with our theoretical framing, neither form of assistance directly altered assortativity; instead, structural reorganization resulted from how hybrid expressive processes reshaped local conversational dynamics.
 Figure 7: Structural equation model of session-level effects of AI assistance on stance assortativity.
@@
 Solid arrows indicate standardized path coefficients, and dashed lines indicate covariances among mediators.
-Black lines indicate paths siginificant at $p<0.001$, and gray lines indicate paths that do not meet this threshold ($N=60$).
+Black lines indicate paths siginificant at $p<0.001$, and gray lines indicate paths that do not meet this threshold.
 Covariance estimates and indirect effects are provided in Extended Data Table 7.
-3 Discussion
-Our experiment demonstrates how LLM-powered message suggestions produce divergent effects on group composition dynamics in online discussions.
-Individual assistance, which tailored suggestions to users’ own stances, increased communication volume and group size but widened stance gaps across groups.
-Relational assistance, which incorporated awareness of other group members’ perspectives, fostered more emotionally positive and receptive exchanges, encouraging participants to form groups with greater opinion diversity.
-These contrasting pathways illustrate how subtle design choices in AI assistance can either entrench division or actively cultivate cohesion at scale.
-Addressing social division and polarization is widely recognized as a pressing societal challenge [16, 1, 3, 29].
-Prior work has tended to treat individual-level and system-level interventions separately [1, 41] with a limited understanding of how effects at one level propagate to the other.
-At the individual level, studies frequently focus on dyadic communication and examine how people might be guided toward consensus on controversial topics [42, 43, 27].
-At the system level, interventions aim to increase opportunities for diverse populations to interact, such as by designing platforms that encourage cross-cutting exposure [32, 44].
-However, these approaches remain disconnected: Individual-level studies lack mechanisms to evolve or even sustain opportunities for ongoing interaction [3], while system-level interventions often fall short because mere exposure is insufficient for social cohesion and can even backfire [32].
-Our work complements these studies by explicitly addressing the dynamics between individual behavior and collective structure [6].
-Rather than aiming for immediate consensus at the interpersonal level [29], we examine how the form of conversation can shape group communication norms and foster connections across differences [45].
-Cohesion at the system level emerged through spontaneous interaction among individuals exercising free choice [46, 47].
-Prior empirical work shows that certain linguistic patterns facilitate the formation of echo chambers online [48, 49].
-Our findings suggest that targeted linguistic interventions can reverse this tendency, counteracting polarization dynamics and fostering more heterogeneous and cohesive group structures.
-This study also highlights the risks and potential of AI assistance in shaping patterns of social division and cohesion.
-As AI technologies continue to advance, they have become embedded in everyday communication tools and are often personalized to individual users [21, 22].
-While such personalization can improve usability, sycophantic AI assistance that echoes users’ views [26] may also amplify perceptual asymmetries, where “in-group virtues become out-group vices” [50].
-This conventional personalization paradigm reflects methodological individualism in social science, which treats individuals as isolated units of analysis [51].
-Our findings suggest complementing this perspective with a relational lens [52].
-Individuals are not fixed entities but enact different roles and identities across social contexts [53, 54, 55].
-If we view a “person” not only as an isolated individual but as someone embedded in interactional networks [56], AI personalization can be designed to help users express their opinions in ways that remain open and responsive to others—even those who disagree [57].
-Human agency is another key consideration in the use of AI assistance.
-In our study, participants sent more messages when suggestions were available but produced fewer self-written ones, and they rarely edited the suggested content.
-These patterns suggest that participants delegated part of their agency to AI, primarily to reduce the effort of composing messages themselves [24].
-As a result, relational assistance facilitated more receptive conversations and greater social cohesion, but these outcomes could occur without participants’ explicit intention.
-This dynamic resembles classic solutions to collective action problems, in which individual incentives help achieve public goods [58].
-Our findings further suggest that humans may learn from AI by adopting more receptive language, as self-written messages became more emotionally positive when exposed to relational suggestions [27].
-At the same time, however, the results raise concerns that human agency and interpersonal trust may erode as control over social interactions shifts toward automated systems [22, 59, 60, 61].
-We acknowledge several limitations of this study that point to directions for future work.
-First, our experiment took place in a controlled online setting with structured discussions, whereas real-world conversations unfold over longer time horizons across multiple platforms [62].
-Participants in our study rarely changed their opinion stance within a brief session, yet over extended interactions, established clusters may further polarize or, conversely, moderate through repeated exposure and trust-building [63, 64].
-Second, future work can also advance the design space of conversational assistance itself.
-Although relational assistance effectively reduced stance-based assortativity, it was used less frequently than individual assistance, which led to higher conversational engagement.
-Exploring additional design dimensions—such as embodiment and controllability—may help balance social receptiveness with engagement [65, 66].
-Finally, beyond technical refinement, evaluating these systems in politically salient contexts will be crucial for understanding how AI assistance operates amid the social and emotional complexities of real-world discourse [17, 15].
-Taken together, this study demonstrates how subtle, AI-mediated changes in communication can scale from individual interactions to collective structures.
-By tracing the micro–macro loop, we show that message-level interventions can cascade into group-composition dynamics, either reinforcing divisions or fostering cohesion in opinion-diverse populations.
-As AI tools become increasingly ubiquitous, they offer an opportunity to redirect their influence toward the production of public goods.
-Reimagining personalization and autonomy as socially embedded processes may help resist the drift toward fragmentation and polarization that increasingly characterizes many societies today.
-Our findings suggest design levers for AI systems and communication policies that could facilitate more cohesive and resilient public discourse.
-4 Methods
-4.1 Ethics, Consent, and Preregistration
+Discussion
+Subtle differences in AI-mediated expressive framing can propagate through conversational dynamics to reorganize the structure of online groups.
+Although assortativity is a session-level measure that usually shifts gradually, we observe clear divergence across conditions.
+Individual assistance, which tailored suggestions to users’ own stances, amplified communication volume yet increased separation between groups.
+Relational assistance, which incorporated awareness of others’ perspectives, fostered more receptive conversations and produced more heterogeneous, cross-cutting group configurations.
+These results indicate that AI-generated phrasing has become part of the interactional substrate through which social structure is produced.
+Addressing the dynamics of social division and polarization remains a critical scientific and societal challenge [42, 15, 16, 30].
+Research has largely bifurcated into two levels of analysis [15, 3].
+Individual-level interventions aim to improve interpersonal receptiveness [43, 44, 27], whereas system-level interventions seek to modify exposure and contact patterns [33, 45].
+Yet how effects at one level translate to the other remains challenging: interpersonal interventions seldom show how conversational shifts propagate into structural outcomes [16], and system-level interventions often falter because externally imposed exposure does not reliably sustain constructive interaction and can even exacerbate division [33].
+Our study links these levels by showing how local expressive cues introduced during conversation can scale into collective structural patterns.
+Rather than targeting immediate persuasion [46, 47] or consensus [30], we focus on how the form of conversation—whether self-focused or socially aware—shapes the interactional norms that govern group formation and reorganization [48].
+This complements prior work showing that specific linguistic styles can intensify ideological clustering [10, 11].
+Our findings demonstrate the converse: when expressive framing highlights relational context, micro-level dynamics can instead dampen segregation pressures.
+The results also illuminate properties of hybrid human–AI expressive systems [20, 22].
+Personalization algorithms tend to mirror users’ existing stances [25, 26], potentially reinforcing perceptual asymmetries in which “in-group virtues become out-group vices” [49].
+This paradigm reflects a form of methodological individualism [50], in which individuals are operationally treated in analytic isolation.
+Our study points to a complementary relational perspective on AI framing [51, 52].
+Individuals enact different roles, identities, and expectations across social contexts [53, 54, 55] and are situated within interactional networks [56].
+Even modest changes in expressive framing—from “help me say what I think” to “help me say what I think in relation to them”—can shift how people engage and how groups subsequently configure themselves.
+Such AI framing can influence group dynamics even when only some participants make use of the suggestions, consistent with evidence that visible expressive cues can shape broader interactional norms [57, 58].
+Human agency is another important consideration.
+In our study, participants sent more messages when suggestions were available but rarely edited them, suggesting partial delegation of expressive effort to the AI tool [24].
+As a result, relational assistance yielded more cohesive structures, yet these outcomes did not require explicit intentions from participants.
+This dynamic resembles classic solutions to collective action problems, where selective incentives align individual behavior with group-level outcomes [59].
+We also find evidence of expressive alignment through dual exposure effects: participants’ self-written messages became more emotionally positive and respectful when they were intermittently exposed both to relational AI suggestions and to sentiment-enhanced messages from group members [27].
+This illustrates that hybrid expressive systems can shape not only message output but also users’ own linguistic patterns through social reinforcement.
+At the same time, increased reliance on automated support, even relational assistance, raises questions about how agency and interpersonal trust may shift as AI mediates a larger fraction of interactional work, and whether stylistic positivity or accommodation might mask substantive disagreement in more consequential settings [19, 60, 61, 62].
+Several limitations suggest directions for future research.
+First, our experiment was conducted in a controlled online environment with short discussions, whereas real interactions unfold over longer time periods and across multiple platforms [63].
+We interpret our results as revealing transferable interactional mechanisms rather than making direct predictions about full-scale online platforms.
+These mechanisms—shifts in conversational tone, asymmetries in message production, and the diffusion of hybrid human–AI expression—could plausibly operate in larger, persistent online communities, making this an important direction for future research.
+For example, in our study, participants rarely changed their own stance, but their opinions may polarize or moderate through longer, repeated exposure and relationship dynamics [64, 65].
+Second, the design space of conversational assistance remains large.
+Although relational assistance reduced assortativity, it was used less frequently than individual assistance.
+Exploring additional forms of expressive support—such as timing, controllability, embodiment, or intent framing—may clarify how assistance can function more effectively as a socio-cognitive scaffold [47, 66, 67, 68].
+Finally, deploying such systems in politically salient or emotionally charged contexts may reveal additional dynamics related to identity, conflict, and misinformation [69, 70].
+This study provides an experimental test of the micro-macro consequences of AI-mediated expression.
+We show that small expressive shifts introduced by AI at the level of individual messages can propagate through conversational interaction to reshape collective outcomes.
+When AI assistance incorporates the social context surrounding users, it fosters more receptive communication and consequently reshapes social structure in more heterogeneous and cohesive ways within opinion-diverse populations.
+As AI systems become increasingly integrated into everyday communication, understanding how hybrid expressive processes diffuse through interaction offers a foundation for explaining—and anticipating—how local expressive cues reshape the macro-level organization of social systems.
+Methods
+Ethics, Consent, and Preregistration
 This study was preregistered [31] and approved by the Carnegie Mellon University Institutional Review Board.
@@
 No identifying information was collected.
-4.2 Participants
+Participants
 We conducted the online experiment from July to August 2025 with 557 unique participants recruited from Prolific [32] across 60 sessions.
@@
 The full task instructions are provided in the Supplementary Information.
-4.3 Procedure
+Procedure
 Upon entry, participants completed a short tutorial and a pre-task survey with seven items, each asking for agreement with a political statement on a 7-point Likert scale (Extended Data Table 1).
@@
 They then received a base payment of \$5.00 for completing the study, with an additional \$1.50 performance bonus contingent on active engagement in the conversation.
-4.4 Message Suggestions
+Message Suggestions
 In the AI assistance conditions, we used OpenAI’s GPT-4o model to generate real-time message suggestions during the conversation.
@@
 Extended Data Table 2 illustrates differences in suggested messages across conditions resulting from this prompt tuning.
-4.5 Analyses
-4.5.1 Group Composition
+Analyses
+Group Composition
 To analyze group composition dynamics, we constructed a bipartite network representation of each session round and projected it onto a participant–participant network, where an edge existed if two participants were in the same group (Fig.
@@
 To enable clearer comparisons, we examined structural measures from Round 2, subtracting the first-round baseline.
-4.5.2 Conversation Dynamics
+Conversation Dynamics
 To analyze conversational patterns within groups, we focused on participants’ message-level behaviors. We measured individual-level messaging activity and the use of AI-generated suggestions.
@@
 4cd).
-4.5.3 Statistical Modeling from Conversation to Cohesion
+Statistical Modeling from Conversation to Cohesion
 To examine how conversational behaviors predicted subsequent ego-centric stance distance within groups, we estimated linear mixed models at the participant–round level.